OpenStudy (anonymous):

What is the axis of symmetry of the function P(n) P(n) = -250n^2 + 3,250n - 9,000

OpenStudy (anonymous):

@ninjasandtigers

myininaya (myininaya):

If you can find the vertex of the parabola, you have done 99.9% of the work.

myininaya (myininaya):

Do you know how to do that?

OpenStudy (anonymous):

nope

myininaya (myininaya):

What do you know about parabolas?

myininaya (myininaya):

Do you know the vertex form of a parabola?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

y=a(x-h)^2+k

myininaya (myininaya):

If I give you the function f(x)=a(x-h)^2+k, would you know how to determine the vertex? This is called the vertex form because it gives you the vertex. What is the vertex for f(x)=a(x-h)^2+k?

OpenStudy (anonymous):

i have no clue how to do that

myininaya (myininaya):

We can use the things we learned about transformations and the graph of y=x^2 to determine that the vertex of f(x)=a(x-h)^2+k is (h,k)

OpenStudy (anonymous):

4 and 9??

myininaya (myininaya):

how did you get that?

OpenStudy (anonymous):

i completed the square

myininaya (myininaya):

we need to write P(n)=-250n^2 + 3,250n - 9,000 in the following form P(n)=a(n-h)^2+k where (h,k) will be the vertex. I will give you the recipe to follow. First pretend we have P(n)=ax^2+bx+c the first thing we want to do is factor a from the first two terms the first terms being ax^2 and bx Now I know you don't see in a in bx but you could write bx as abx/a since a/a is 1 and abx/a is still bx so anyways P(n)=ax^2+abx/a+c Factor out out a from the first two terms like so P(n)=a(x^2+bx/a)+c <--do this as a first step with your function then I will give the next step

OpenStudy (anonymous):

whattt im so confused i dont even know what this is

myininaya (myininaya):

what is?

myininaya (myininaya):

oops I put x's instead of n's P(n)=an^2+abn/a+c P(n)=a(n^2+bn/a)+c

myininaya (myininaya):

All this is saying I want you to look at the coeifficent of n^2 which your a here is -250 and factor out your a from the first two terms that is all the first step requires

OpenStudy (anonymous):

vertex= 6.5 and 1562.5

OpenStudy (anonymous):

i found the vertex

myininaya (myininaya):

well the vertex would be in this form (h,k) I guess you are saying h is 6.5 and k is 1562.5 bascially the vertex is the center of the graph or what i mean is it is where you can draw a line and see symmetry I will give a rough picture of it  You are suppose to be able to fold your graph about that vertical line I drew and it is suppose to lay perfectly on the other side. Do you see how that vertical line goes through the vertex? If you know the x-coordinate of the vertex and you know that vertical lines are x=constant then you should be able to come up with the axis of symmetry (the line that divides the graph into two equal halves)